OpenStudy (anonymous):

What is equivalent to 4(2x+1)-(-6x)

OpenStudy (mathmale):

1. Multiply out 4(2x+1). 2. Multiply out -(-6x). 3. Combine the results. I deleted your other post because it was the same as this one. A warm welcome to OpenStudy!

OpenStudy (jmark):

Multiply the terms 4(2x+1)-(-6x) 8x+4- (-6x) = 0 8x+4+6x =0 14x+4=0 14x = -4 x = -4/14 x = -2/7 ICSE syllabus class 6 - http://icse.edurite.com/icse-syllabus/icse-syllabus-for-class-6.html

OpenStudy (mathmale):

@Neilvel: As before, i hope your experience on OpenStudy turns out to be very positive. Note, however, that to make the most of this online rsource, you'll need to get involve and to respond when someone tries to help you.